THE MICROWAVE SPECTRUM OF 1,2,4-TRIAZINE AND THE ROTATIONAL CONSTANTS OBTAINED FROM A SIMULTANEOUS ANALYSIS OF MICROWAVE GROUND STATE AND HIGH RESOLUTION IR-TRANSITIONS.

\underline{DINES CHRISTEN}, Institute of Physical and Theoretical Chemistry, University of Tübingen, Germany; MICHAEL H. PALMER, Department of Chemistry, University of Edinburgh, Scotland, UK; FLEMMING HEGELUND, Department of Chemistry, Aarhus University, Denmark.

The microwave spectrum of 1,2,4-triazine has been recorded in the Ku- and Ka-bands using a conventional stark spectrometer. The low-J lines appeared close to the frequencies predicted for the ground state from the high resolution FTIR-studies^{a}, but to secure a correct assignment of higher J Q-branch transitions, a microwave-microwave double resonance experiment was done for a few pairs of transitions sharing a common level. 145 transitions were recorded with J_{max}= 42, K_{a,max}= 33, and K_{c,max}= 11. A simultaneous fit to MW pure rotational transitions and to the rotational structure of selected IR-bands led to a determination of rotational and all 5 quartic centrifugal distortion constants.

aPalmer, Maier, Hegelund, and Newnham, J. Mol. Spectrossc. 192, 331, (1998) and Bach, Hegelund, Beukes, Nicolaisen, and Palmer, J. Mol. Spectrosc. 198, (1999)